# Holographic dictionary for Lifshitz and hyperscaling violating black holes

### v1

@@ Abstract
Introduction. Holography, or gauge/gravity duality, is the idea that a theory of quantum gravity in a ($d+1$)-dimensional bulk spacetime is dual to a quantum theory without gravity living on the $d$-dimensional asymptotic boundary of the bulk spacetime [1, 2]. This duality is best understood for gravitational theories in asymptotically Anti-de Sitter (AdS) spacetime, for which the dual theory is a conformal field theory (CFT), also known as the ‘AdS/CFT correspondence’ [3, 4, 5]. An important feature of this correspondence is that the thermodynamics of stationary AdS black holes [6] is dual to the thermodynamics of high-energy equilibrium states in a strongly coupled, large-$N$ conformal gauge theory [7]. For instance, the black hole entropy $S=A/4G$ [8, 9] is identified with the thermodynamic entropy $\tilde{S}$ of the CFT equilibrium state, and the internal energy $\tilde{E}$, temperature $\tilde{T}$, electric potential $\tilde{\Phi}$ and charge $\tilde{Q}$ of the CFT state are proportional to, respectively, the mass $M$, temperature $T$, electric potential $\Phi$ and charge $Q$ of the black hole [10, 11, 12, 13]
@@
The general principles behind holography are expected to apply to non-asymptotically AdS spacetimes as well. Since the isometries of the bulk spacetime should match with the global symmetries of the dual boundary theory, field theories that are not conformally (or scale) invariant are not dual to AdS spacetime, but possibly to a different bulk geometry. For example, some quantum field theories exhibit the anisotropic scaling symmetry $\{t,x^{i}\}\to\{\alpha^{z}t,\alpha x^{i}\}$, parametrized by the dynamical critical exponent $z$. Together with time and space translations and space rotations, this scaling symmetry forms the Lifshitz symmetries. For $z\neq 1$, Lifshitz invariant theories describe non-relativistic quantum critical systems. The anisotropic scaling symmetry can be geometrically realized in the bulk by the so-called Lifshitz spacetimes, which are a generalization of AdS spacetime [15, 16, 17, 18] (see [19] for a review). Another generalization is given by hyperscaling violating geometries, which are dual to theories violating the hyperscaling property that entropy (or free energy) scales with its naive dimension. For hyperscaling violating Lifshitz (HVL) theories (on the plane), entropy scales with temperature as $S\propto T^{(d-1-\theta)/z}$, where $\theta$ is the hyperscaling violation exponent [20, 21]. Various black hole solutions with Lifshitz and/or hyperscaling violating asymptotics have been obtained in [18, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33].
@@
$\tilde{E}=\tilde{T}\tilde{S}+\tilde{\Phi}\tilde{Q}+\mu C\,,$ (2)
where the central charge $C$ is defined as the dimensionless proportionality factor of the grand canonical free energy, and $\mu$ is its conjugate chemical potential (we suppress the tildes on $\mu,C$ since they do not have an equivalent definition in the bulk). This Euler relation is valid for any large-$C$ field theory at finite temperature in the deconfined phase and it is the same for theories with different global symmetries (e.g., conformal and Lifshitz). Moreover, we show that an extended version of the first law for these black holes is dual to the thermodynamic first law on the boundary
@@
Lifshitz and hyperscaling violating black holes. A line element of the general form [^1] Note that we are working with $d+1$ bulk spacetime dimensions, whereas [33] has $d+2$ bulk dimensions. Moreover, we redefined $V_{0}$ in [33] as $-2\Lambda_{0}$, and $\phi_{0}$ as $-\gamma\log r_{0}$, see (11).
@@
in $(d+1)$-dimensions, with $f(r)\rightarrow 1$ as $r\rightarrow\infty$, has been suggested to be dual to states in Lifshitz invariant field theories with hyperscaling violation. For $z=1$ and $\theta=0$ we recover asymptotically AdS spacetime. Here, $k$ parameterizes the topology of the $(t,r)=$ constant surfaces: $k=\{-1,0,1\}$ for hyperbolic, planar and spherical topology respectively. When $f(r)=1$ under the rescalings
@@
with $\alpha$ being a dimensionless scaling parameter. Leaving aside the scaling of the bulk radius $r$, this rescaling of the boundary metric is characteristic of hyperscaling violating field theories. Hence, we refer to (4) as asymptotically Lifshitz and hyperscaling violating geometries (although, strictly speaking, this is only true for $k=0$, since geometries with $k=\pm 1$ are not invariant under spatial translations.) As is standard for hyperscaling violating theories, the dual black hole geometry is considered to be valid up to a certain large radius $r_{F}$, which corresponds to a UV cut-off scale on the boundary, since the dual field theory could have a different UV completion [28, 31, 34].
[…]
@@
where the $\lambda_{i}$ are given in terms of ($d,\theta,z)$; namely:
@@
allowing for a non-zero hyperscaling violating parameter $\theta$. The gauge field A is responsible for the anisotropic scaling symmetry, B supports non-trivial (spherical or hyperbolic) horizon topology, and C allows for non-zero electric charge. The analytic solution for these black holes is given by (4) with [33]
@@ C $=-\frac{\rho_{3}}{d-\theta+z-3}r_{0}^{\gamma\lambda_{3}}r^{-(d-\theta+z-3)}dt\,,$ (10)
generalizing previously discovered solutions for charged spherical black holes with Lifshitz asymptotics [27] (to non-zero $\theta$) and charged black branes with arbitrary Lifshitz and hyperscaling violation parameters [29] (to non-trivial topology). The respective mass and charge parameters are $m$ and $q$, and
@@
Extended bulk thermodynamics. The thermodynamic quantities of the hyperscaling violating Lifshitz black hole solution were computed in [33] (see also, e.g., [27, 29, 35, 36, 37]). The Hawking temperature $T$, Bekenstein-Hawking entropy $S$, ADM mass $M$, electric charge $Q$, and electric potential $\Phi$ are [^2] The ADM mass was obtained in [33] through the background subtraction method, so that the solution with $m=0$ has zero total gravitational energy. For hyperbolic black holes, however, this solution does not correspond to the ground state of the grand canonical ensemble, which is given by the extremal solution with zero charge and in our conventions hence has negative energy (whereas in [33] the ADM mass for $k=-1$ was redefined such that the extremal solution has zero energy).
@@
From the mass function (35) the following scaling relation may be inferred:
@@
the conjugate quantity to $P$. If $\theta=0$ and $z=1$, $\Theta$ reduces to the expected expression $-r^{d}\omega_{k,d-1}/d$. Hence $\Theta$ (see equation (42)) is the HVL generalization of (minus) the “thermodynamic volume”, see e.g. [39].
The generalized Smarr relation (16) is new in the literature, as the scaling relation (15) was not properly derived before. In fact, in [35, 33, 40] it was assumed that the standard Smarr formula for AdS black holes (given by (16) with $z=1$ and $\theta=0$) also holds HVL black holes. However this is clearly not the case, since the Smarr relation (16) explicitly involves $z$ and $\theta$.
@@
The dictionary for the central charge can be read off from the proportionality constant of the grand canonical free energy $F=M-TS-\Phi Q$. Since the mass, entropy and charge all scale with the central charge, $M,S,Q\propto C$, we can directly determine its dictionary entry by looking at, for instance, the Bekenstein-Hawking entropy (Holographic dictionary for Lifshitz and hyperscaling violating black holes). The central charge is dimensionless and should only depend on theory parameters, not on the specific details of the state, such as its temperature. Therefore, the dictionary for $C$ cannot involve the horizon radius $r_{h}$, which is related to the temperature (Holographic dictionary for Lifshitz and hyperscaling violating black holes). Multiplying and dividing the entropy by $L^{d-\theta-1}$, we can write
@@
with $x=r_{h}/L$, where we identify the central charge as
$C\propto\frac{L^{d-\theta-1}r_{F}^{\theta}}{G}\,.$ (23)
For $\theta=0$ this dictionary reduces to the well-known result for Einstein gravity. It is interesting to note that the dictionary entry (23) depends on $\theta$ but not $z$. It is well known that $\theta$ effectively lowers the number of dimensions, which is also apparent from (23). Moreover, it is quite natural that (23) involves $r_{F}$, since that corresponds to a UV cut-off in the boundary theory, on which the number of field degrees of freedom depends. Likewise, the entropy does not depend on $r_{0}$ and so the entry (23) for $C$ does not, since $S\propto C$.
The dictionary for the internal energy and temperature follows from symmetry arguments. The products $ML$ and $TL$ in the bulk are dimensionless, and so should match with products of boundary quantities that are (from (5)) invariant under the anisotropic scaling symmetry, $\{t,x^{i}\}\to\{\alpha^{z}t,\alpha x^{i}\}$, and the simultaneous hyperscaling transformation $\{t,x^{i}\}\to\alpha^{-\theta/(d-1)}\{t,x^{i}\}$ that undoes the rescaling (6) of the metric at the boundary. The invariant combinations are
@@
Note that the scaling dimension of energy and temperature is $[\tilde{E}]=[\tilde{T}]=-[t]=z-\theta/(d-1)$, and the scaling dimension of the boundary curvature scale and spatial coordinate is $[R]=[x^{i}]=-1+\theta/(d-1)$ (see Appendix D in [36]), so that the LHS of (24) is indeed dimensionless.
The dictionary entries for the electric potential and charge are more complicated, but also essentially follow from dimensional analysis. We take the boundary electric charge $\tilde{Q}$ to be dimensionless, as for the entropy and central charge in (20); the boundary potential $\tilde{\Phi}$ then has the same dimensions as $\tilde{E}$ and $\tilde{T}$, so that $\tilde{\Phi}R^{\frac{z-\theta/(d-1)}{1-\theta/(d-1)}}$ is scale invariant. The bulk charge scales as $Q\propto qL^{-1}r_{F}^{\theta-\theta/(d-1)}r_{0}^{1-z+\theta/(d-1)}/G$ and the bulk potential as $\Phi\propto qr_{F}^{\theta/(d-1)}r_{0}^{z-1-\theta/(d-1)}/(L^{z}r_{h}^{d-\theta+z-3})$, where the charge parameter and Newton’s constant have mass dimension $[q]=-d+\theta-z+2$ and $[G]=-d+1$, respectively. We can thus make the boundary charge $\tilde{Q}$ and the product $\tilde{\Phi}R^{\frac{z-\theta/(d-1)}{1-\theta/(d-1)}}$ dimensionless with the following dictionary
@@
Note this product is independent of $r_{0}$ and $r_{F}$. For $z=1$ and $\theta=0$ this agrees with the standard holographic dictionary [11]. There are no $\sqrt{G}$ factors due to a different normalization of the action compared to that in [13] (in our conventions the gauge coupling constant depends on $G$). We further comment that this corrects an earlier dictionary (Appendix D in [11]) for the electric charge and potential of Lifshitz black holes, where it was proposed that $\tilde{\Phi}=\Phi/R^{z}$ and $\tilde{Q}=QL$, which is different from the dictionary above for $\theta=0$, i.e. $\tilde{\Phi}=\Phi L^{z}/R^{z}$ and $\tilde{Q}=QL^{2-z}$. This difference ultimately arises from the fact that in [11] (following [35, 33]) it was assumed that the AdS Smarr formula also holds for Lifshitz black holes, whereas here we find that a more careful study of the scaling relation for the mass shows that the correct Smarr formula is (16).
Crucially, using the dictionary in (21) and (23)-(26) one can show that the extended bulk first law (19) exactly matches with the boundary first law (20). [^3] We also checked that the matching between the bulk and boundary extended first laws continues to hold if $r_{F}$ and $r_{0}$ are allowed to vary in the bulk (see Supplemental Material). The matching only works if the pressure satisfies the equation of state for HVL theories [19, 36]
$p=\frac{\tilde{E}}{V}\frac{z-\theta/(d-1)}{d-1-\theta}\,,$ (28)
and if the chemical potential $\mu$ obeys the holographic Euler equation (2). The equation of state (28) is a consequence of the scaling relation
which follows from the anisotropic scaling symmetry and the hyperscaling violating property. Differentiating (29) with respect to $\alpha$ and setting $\alpha=1$ yields (28), where the pressure is defined as $p:=-(\partial\tilde{E}/\partial V)_{\tilde{S},\tilde{Q},C}$.
On a planar boundary geometry (for $k=0$) there is an additional relation between the thermodynamic quantities, which is dual to an extra Smarr formula in the bulk. This allows us to replace the $\mu C$ term in the Euler relation with a $pV$ term, but this comes at the expense of additional $(z,\theta,d)$-dependent prefactors in front of the $pV$ and $\tilde{\Phi}\tilde{Q}$ terms (that are absent for $z=1,\theta=0$), namely:
@@
We refer to the Supplemental Material for more details.
@@
$2\Theta P$ $=R^{\frac{z-\theta/(d-1)}{1-\theta/(d-1)}}\frac{1}{L}\left((2-z)\tilde{\Phi}\tilde{Q}+(d-\theta-1)\mu C-\tilde{E}\right)\,.$ (34)
@@
For future work, it would be interesting to derive the dictionary for the thermodynamic black hole quantities using holographic renormalization [42, 43, 44] and, in particular, check our formula for the central charge (which can be read off from the Casimir energy for even $d$ and $k\neq 0$). Moreover, we would like to understand whether there exists a thermodynamic interpretation of the length parameters $r_{F}$ and $r_{0}$. Finally, it is likely that our new setup will lead to a discovery of so far unseen phase transitions and critical phenomena arising for Lifshitz and hyperscaling violating theories.
Acknowledgments. D.K. is grateful for support from GAČR 23-07457S grant of the Czech Science Foundation and the Charles University Research Center Grant No. UNCE24/SCI/016. M.R.V. acknowledges support from the SNF Postdoc Mobility grant P500PT-206877 and the Spinoza Grant of the Dutch Science Organisation (NWO). R.B.M. acknowledges support from the Natural Sciences and Engineering Research Council of Canada.

### v2

@@ Abstract
Introduction. Holography, or gauge/gravity duality, is the idea that a theory of quantum gravity in a ($d+1$)-dimensional bulk spacetime is dual to a quantum theory without gravity living on the $d$-dimensional asymptotic boundary of the bulk spacetime [1, 2]. This duality is best understood for gravitational theories in asymptotically Anti-de Sitter (AdS) spacetime, for which the dual theory is a conformal field theory (CFT), also known as the ‘AdS/CFT correspondence’ [3, 4, 5]. An important feature of this correspondence is that the thermodynamics of stationary AdS black holes [6] is dual to the thermodynamics of high-energy equilibrium states in a strongly coupled, large-$N$ conformal gauge theory [7]. For instance, the black hole entropy $S=A/4G$ [8, 9] is identified with the thermodynamic entropy $\tilde{S}$ of the CFT equilibrium state, and the internal energy $\tilde{E}$, temperature $\tilde{T}$, electric potential $\tilde{\Phi}$, and charge $\tilde{Q}$ of the CFT state are proportional to, respectively, the mass $M$, temperature $T$, electric potential $\Phi$ and charge $Q$ of the black hole [10, 11, 12, 13]
@@
The general principles behind holography are expected to apply to nonasymptotically AdS spacetimes as well. Since the isometries of the bulk spacetime should match with the global symmetries of the dual boundary theory, field theories that are not conformally (or scale) invariant are not dual to AdS spacetime, but possibly to a different bulk geometry. For example, some quantum field theories exhibit the anisotropic scaling symmetry $\{t,x^{i}\}\to\{\alpha^{z}t,\alpha x^{i}\}$, parametrized by the dynamical critical exponent $z$. Together with time and space translations and space rotations, this scaling symmetry forms the Lifshitz symmetries. For $z\neq 1$, Lifshitz invariant theories describe nonrelativistic quantum critical systems. The anisotropic scaling symmetry can be geometrically realized in the bulk by the so-called Lifshitz spacetimes, which are a generalization of AdS spacetime [15, 16, 17, 18] (see [19] for a review). Another generalization is given by hyperscaling violating geometries, which are dual to theories violating the hyperscaling property that entropy (or free energy) scales with its naive dimension. For hyperscaling violating Lifshitz (HVL) theories (on the plane), entropy scales with temperature as $S\propto T^{(d-1-\theta)/z}$, where $\theta$ is the hyperscaling violation exponent [20, 21]. Various black hole solutions with Lifshitz and/or hyperscaling violating asymptotics have been obtained in [18, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33].
@@
$\tilde{E}=\tilde{T}\tilde{S}+\tilde{\Phi}\tilde{Q}+\mu C\,.$ (2)
The central charge $C$ is defined as the dimensionless proportionality factor of the energy, entropy and electric charge in the canonical ensemble, i.e., $\tilde{\chi}=Cf_{\tilde{\chi}}(\tilde{T},\tilde{Q},V)$ for $\tilde{\chi}\in\{\tilde{E},\tilde{S},\tilde{Q}\}$. Furthermore, the chemical potential is defined as the conjugate quantity to $C$ in the first law, as $\mu\equiv(\partial\tilde{E}/\partial C)_{\tilde{S},\tilde{Q},V}.$ We suppress the tildes on $\mu,C$ since they do not have an equivalent definition in the bulk. The Euler relation (2) follows from the nonextensive scaling relation $\tilde{E}(\alpha\tilde{S},V,\alpha\tilde{Q},\alpha C)=\alpha\tilde{E}(\tilde{S},V,\tilde{Q},C)$, where $V$ is the spatial volume. This scaling relation only holds in the large-$C$ limit, since there are $1/C$ corrections at finite central charge. More specifically, the Euler relation is valid for any large-$C$ field theory at finite temperature in the deconfined phase. Moreover, it is the same for theories with different global symmetries, e.g., conformal and Lifshitz symmetries.
Furthermore, we show that an extended version of the first law for these black holes is dual to the thermodynamic first law on the boundary
@@
Lifshitz and hyperscaling violating black holes. A line element of the general form [^1] Note that we are working with $d+1$ bulk spacetime dimensions, whereas [33] has $d+2$ bulk dimensions. Moreover, we redefined $V_{0}$ in [33] as $2\Lambda_{0}$, and $\phi_{0}$ as $\gamma\log r_{0}$, see (11).
@@
in $(d+1)$ dimensions, with $f(r)\rightarrow 1$ as $r\rightarrow\infty$, has been suggested to be dual to states in Lifshitz invariant field theories with hyperscaling violation. For $z=1$ and $\theta=0$ we recover asymptotically AdS spacetime. Here, $k$ parametrizes the topology of the $(t,r)=$ constant surfaces: $k=\{-1,0,1\}$ for hyperbolic, planar and spherical topology respectively. When $f(r)=1$ under the rescalings
@@
with $\alpha$ being a dimensionless scaling parameter. Leaving aside the scaling of the bulk radius $r$, this rescaling of the boundary metric is characteristic of hyperscaling violating field theories. Hence, we refer to (4) as asymptotically Lifshitz and hyperscaling violating geometries (although, strictly speaking, this is only true for $k=0$, since geometries with $k=\pm 1$ are not invariant under spatial translations). As is standard for hyperscaling violating theories, the dual black hole geometry is considered to be valid up to a certain large radius $r_{F}$, which corresponds to a UV cutoff scale on the boundary, since the dual field theory could have a different UV completion [28, 31, 34].
Static, charged black holes with line element (4) were found as analytic solutions to a generalized Einstein-Maxwell-Dilaton theory [33], consisting of Einstein gravity minimally coupled to a real scalar field $\phi$, called the “dilaton,” and three Abelian gauge fields $\textsf{A},\textsf{B}$ and C. The action of the generalized EMD theory is given by
@@
where the $\lambda_{i}$ are given in terms of ($d,\theta,z)$; namely,
@@
allowing for a nonzero hyperscaling violating parameter $\theta$. The gauge field A is responsible for the anisotropic scaling symmetry, B supports nontrivial (spherical or hyperbolic) horizon topology, and C allows for nonzero electric charge. The analytic solution for these black holes is given by (4) with [33]
@@ C $=-\frac{\rho_{3}}{d-\theta+z-3}r_{0}^{\gamma\lambda_{3}}r^{-(d-\theta+z-3)}dt\,$, (10)
generalizing previously discovered solutions for charged spherical black holes with Lifshitz asymptotics [27] (to nonzero $\theta$) and charged black branes with arbitrary Lifshitz and hyperscaling violation parameters [29] (to nontrivial topology). The respective mass and charge parameters are $m$ and $q$, and
@@
Extended bulk thermodynamics. The thermodynamic quantities of the hyperscaling violating Lifshitz black hole solution were computed in [33] (see also, e.g., [27, 29, 35, 36, 37]). The Hawking temperature $T$, Bekenstein-Hawking entropy $S$, ADM mass $M$, electric charge $Q$, and electric potential $\Phi$ are [^2] The ADM mass was obtained in [33] through the background subtraction method, so that the solution with $m=0$ has zero total gravitational energy. For hyperbolic black holes, however, this solution does not correspond to the ground state of the grand canonical ensemble, which is given by the extremal solution with zero charge and in our conventions hence has negative energy (whereas in [33] the ADM mass for k=1k=-1 was redefined such that the extremal solution has zero energy).
@@
From the mass function (35) in the Supplemental Matertial the following scaling relation may be inferred:
@@
the conjugate quantity to $P$. If $\theta=0$ and $z=1$, $\Theta$ reduces to the expected expression $r^{d}\omega_{k,d-1}/d$. Hence $\Theta$ (see equation (42) in the Supplemental Matertial) is the HVL generalization of (minus) the “thermodynamic volume”, see e.g. [39].
The generalized Smarr relation (16) is new in the literature, as the scaling relation (15) was not properly derived before. In fact, in [35, 33, 40] it was assumed that the standard Smarr formula for AdS black holes (given by (16) with $z=1$ and $\theta=0$) also holds for HVL black holes. However this is clearly not the case, since the Smarr relation (16) explicitly involves $z$ and $\theta$.
@@
The holographic dictionary follows from requiring (i) that the extended bulk first law (19) exactly matches with the boundary first law (20); (ii) that the chemical potential $\mu$ obeys the holographic Euler equation (2); and (iii) that the pressure $p$ satisfies the equation of state for HVL theories [19, 36]
$p=\frac{\tilde{E}}{V}\frac{z-\theta/(d-1)}{d-1-\theta}\,.$ (22)
These three conditions yield the dictionary given below in (24)-(28). [^3] We also checked that the matching between the bulk and boundary extended first laws continues to hold if $r_{F}$ and $r_{0}$ are allowed to vary in the bulk (see Supplemental Material). We note that this equation of state is a consequence of the scaling relation
which follows from the anisotropic scaling symmetry and the hyperscaling violating property. Differentiating (23) with respect to $\alpha$ and setting $\alpha=1$ yields (22), where the pressure is defined as $p:=-(\partial\tilde{E}/\partial V)_{\tilde{S},\tilde{Q},C}$. We now discuss the dictionary for the various thermodynamic parameters in turn.
First, we find the following dictionary for the central charge
$C\propto\frac{L^{d-\theta-1}r_{F}^{\theta}}{G}\,.$ (24)
For $\theta=0$ this dictionary reduces to the well-known result for Einstein gravity. The central charge is dimensionless and should only depend on theory parameters, not on the specific details of the state, such as its temperature. Hence, the dictionary for $C$ does not involve the horizon radius $r_{h}$, which is related to the temperature (Holographic dictionary for Lifshitz and hyperscaling violating black holes). Further, it is interesting to note that the right-hand side of (24) depends on $\theta$ but not $z$. It is well known that $\theta$ effectively lowers the number of dimensions, which is also apparent from (24). Moreover, it is quite natural that (24) involves $r_{F}$, since that corresponds to a UV cutoff in the boundary theory, on which the number of field degrees of freedom depends.
We note the dictionary (24) is consistent with the definition of $C$ as the proportionality constant of the energy, entropy and electric charge, i.e., $\tilde{E},S,\tilde{Q}\propto C$. For instance, the Bekenstein-Hawking entropy (Holographic dictionary for Lifshitz and hyperscaling violating black holes) is proportional to the central charge,
@@
where we defined $x\equiv r_{h}/L$. Since $S\propto C$ and the entropy does not depend on $r_{0}$, the entry (24) for $C$ likewise does not.
Second, the dictionary for the internal energy and temperature follows partly from symmetry arguments. We can make the internal energy and temperature dimensionless by simply multiplying them with the bulk curvature radius $L$. On the field theory side, this should match with products of boundary quantities that are (from (5)) invariant under the anisotropic scaling symmetry, $\{t,x^{i}\}\to\{\alpha^{z}t,\alpha x^{i}\}$, and the simultaneous hyperscaling transformation $\{t,x^{i}\}\to\alpha^{-\theta/(d-1)}\{t,x^{i}\}$ that undoes the rescaling (6) of the metric at the boundary. The invariant combinations are
@@
Note that the scaling dimension of energy and temperature is $\tilde{E}]=[\tilde{T}]=-[t]=z-\theta/(d-1)$, and the scaling dimension of the boundary curvature scale and spatial coordinate is $R]=[x^{i}]=-1+\theta/(d-1)$ (see Appendix D in [36]), so that the lhs of (26) is indeed dimensionless.
Next, the dictionary entries for the electric potential and charge are more complicated, but also follow for a large part from dimensional analysis. We take the boundary electric charge $\tilde{Q}$ to be dimensionless, as for the entropy and central charge in (20); the boundary potential $\tilde{\Phi}$ then has the same dimensions as $\tilde{E}$ and $\tilde{T}$, so that $\tilde{\Phi}R^{\frac{z-\theta/(d-1)}{1-\theta/(d-1)}}$ is scale invariant. The bulk charge scales as $Q\propto qL^{-1}r_{F}^{\theta-\theta/(d-1)}r_{0}^{1-z+\theta/(d-1)}/G$ and the bulk potential as $\Phi\propto qr_{F}^{\theta/(d-1)}r_{0}^{z-1-\theta/(d-1)}/(L^{z}r_{h}^{d-\theta+z-3})$, where the charge parameter and Newton’s constant have mass dimension $q]=-d+\theta-z+2$ and [G]=d+1[G]=-d+1, respectively. We can thus make the boundary charge $\tilde{Q}$ and the product $\tilde{\Phi}R^{\frac{z-\theta/(d-1)}{1-\theta/(d-1)}}$ dimensionless with the following dictionary
@@
Note this product is independent of $r_{0}$ and $r_{F}$.
For $z=1$ and $\theta=0$ this agrees with the standard holographic dictionary [11]. There are no $\sqrt{G}$ factors due to a different normalization of the action compared to that in [13] (in our conventions the gauge coupling constant depends on $G$). We further comment that this corrects an earlier dictionary (Appendix D in [11]) for the electric charge and potential of Lifshitz black holes, where it was proposed that $\tilde{\Phi}=\Phi/R^{z}$ and $\tilde{Q}=QL$, which is different from the dictionary above for $\theta=0$, i.e. $\tilde{\Phi}=\Phi L^{z}/R^{z}$ and $\tilde{Q}=QL^{2-z}$. This difference ultimately arises from the fact that in [11] (following [35, 33]) it was assumed that the AdS Smarr formula also holds for Lifshitz black holes, whereas here we find that a more careful study of the scaling relation for the mass shows that the correct Smarr formula is (16).
We emphasize that the left-hand side of the holographic dictionary (26)-(28) follows from invariance under Lifshitz and hyperscaling transformations. However, in order to fix the right-hand side of these equations, we imposed that the boundary and bulk extended thermodynamics should be consistent with each other. Specifically, we imposed the three conditions stated above (22).
On a planar boundary geometry (for $k=0$) there is an additional relation between the thermodynamic quantities, which is dual to an extra Smarr formula in the bulk. This allows us to replace the $\mu C$ term in the Euler relation with a $pV$ term, but this comes at the expense of additional $(z,\theta,d)$-dependent prefactors in front of the $pV$ and $\tilde{\Phi}\tilde{Q}$ terms (that are absent for $z=1,\theta=0$), namely,
@@
This implies that the grand canonical free energy $\tilde{F}=\tilde{E}-\tilde{T}\tilde{S}-\tilde{\Phi}\tilde{Q}$ is not equal to pV-pV for generic $z,\theta.$ We refer to the Supplemental Material for more details.
@@
[…]
@@
For future work, it would be interesting to derive the dictionary for the thermodynamic black hole quantities using holographic renormalization [42, 43, 44] and, in particular, check our formula for the central charge (which can be read off from the Casimir energy for even $d$ and $k\neq 0$). Moreover, we would like to understand whether there exists a thermodynamic interpretation of the length parameters $r_{F}$ and $r_{0}$, perhaps relating these to “Lifshitz charges”, as proposed in a simplified setting in [36]. Finally, it is likely that our new setup will lead to a discovery of so far unseen phase transitions and critical phenomena arising for Lifshitz and hyperscaling violating theories.
Acknowledgments. D.K. is grateful for support from Grant No. GAČR 23-07457S of the Czech Science Foundation and the Charles University Research Center Grant No. UNCE24/SCI/016. M.R.V. acknowledges support from the Swiss National Science Foundation Postdoc Mobility Grant No. P500PT-206877 and the Spinoza Grant of the Dutch Science Organisation (NWO). R.B.M. acknowledges support from the Natural Sciences and Engineering Research Council of Canada.